# Radar Signal Recognition through Self-Supervised Learning and Domain Adaptation

### v2

@@ IV-C Ablation Study
We examine the impact of various design considerations for few-shot RSR. As discussed in Section IV-B, the choice of $S_{\text{m}}$ and $R_{\text{m}}$ influences SSL performance and, in turn, fine-tuning and test performance. Although no particular masking strategy consistently outperformed others across our experiments, we found that the masking ratio played a more important role in determining test performance (Table II). While a small batch size coupled with a low learning rate generally benefits fine-tuning, we observed that freezing the model weights for a few epochs provided a slight improvement (less than $5\%$) in test performance. This effect was consistent for models pre-trained on both in-domain and out-of-domain data. We also explored $\ell_{2}$ loss for pre-training, but we observed no meaningful improvements over $\ell_{1}$ loss in our experiments.
V Conclusion
@@
In this paper, we introduced masked signal modelling as an effective self-supervised pre-training method for few-shot radar signal recognition. We also demonstrated the viability of RF domain adaptation for enhancing signal classification performance when no target domain data was used for pre-training. Our results show that by optimally designing the masking method during pre-training, fine-tuned models can achieve significant performance improvements, particularly in moderate to high SNR settings. This is demonstrated by the simple ResNet-based MAE, which achieved a boost in classification accuracy of $17.5\%$ when pre-trained on in-domain signals (i.e., RadChar-SSL) and $16.31\%$ when pre-trained on out-of-domain signals (i.e., RadioML), compared to its $1$-shot baseline without SSL. In future work, additional domains and downstream tasks (e.g., pulse activity segmentation) will be explored to assess the practical utility of SSL and RF domain adaptation for few-shot radar signal recognition.

### v3

@@ 3.3 Ablation Study
We examine the impact of various design considerations for few-shot RSR. As discussed in Section 3.2, the choice of $S_{\text{m}}$ and $R_{\text{m}}$ influences SSL performance and, in turn, fine-tuning and test performance. Although no particular masking strategy consistently outperformed others across our experiments, we found that the masking ratio played a more important role in determining test performance (Table 2). While a small batch size coupled with a low learning rate generally benefits fine-tuning, we observed that freezing the model weights for a few epochs provided a slight improvement (less than $5\%$) in test performance. This effect was consistent for models pre-trained on both in-domain and out-of-domain data. We also explored $\ell_{2}$ loss for pre-training, but we observed no meaningful improvements over $\ell_{1}$ loss in our experiments.
4 Conclusion
@@
In this paper, we introduced MSM as an effective SSL method for few-shot RSR. We also demonstrated the viability of RF domain adaptation for enhancing signal classification performance when no target domain data was used for pre-training. Our results show that by optimally designing the masking method during pre-training, fine-tuned models can achieve significant performance improvements, particularly in moderate to high SNR settings. This is demonstrated by ResNet1D, which achieved a boost in classification accuracy of $17.5\%$ when pre-trained on in-domain signals (RadChar-SSL), and $16.31\%$ when pre-trained on out-of-domain data (RadioML), compared to its $1$-shot baseline without SSL. In future work, additional downstream tasks may be explored to evaluate SSL and RF domain adaptation in a real-world setting.